Shelby works for Dynasty Motors, an automobile dealership. All employees can buy a car at the company's cost plus 5%. The company does not charge employees the $400 dealer preparation fee that nonemployees must pay. Shelby purchased an automobile for $22,785 ($21,700 + $1,085). The company's cost was $21,700. The price for a nonemployee would have been $23,100 ($22,700 + $400 preparation fee).

How much must Shelby include in his gross income related to the purchase of the car?

Tip: The answer is not 320 or 300.
Answer #1

The Discount on the price of the automobile of $1085 (22785-21700) is qualified for Employee Discount

The Gross Profit component of the price to customer is = 22700-21700 = 1000.

Therefore 1085 - 1000 = 85 of the discount is included in Shelby's Gross income.

In addition, the maximum qualified employee discount that can be excluded for a service is 20%

hence Shelby must include 80% of Preparation fee 400 = 320 in Gross income.

Hence the Total taxable amount is 320 + 85 = $ 405.